# MSCN: Multi-view Structural Convolution Network for Domain-Invariant Point Cloud Recognition of Autonomous Vehicles

### v4

@@ V-B Evaluation on Synthetic Datasets: Robustness to Geometric Deformation
MSCN (Ours) 92
V-C Evaluation on Real-World Datasets: Comparable Performance and Robustness
We conduct experiments on the ScanObjectNN [67] dataset to evaluate the performance of MSCN on real-world indoor point clouds. Table III summarizes the classification results, demonstrating that MSCN achieves performance comparable to the SOTA models in standard evaluation settings. Furthermore, MSCN exhibits remarkable robustness under geometric transformations such as scaling and translation, maintaining stable performance and demonstrating its strength in handling these variations.
However, it is important to note that ScanObjectNN consists of point clouds captured in indoor environments, which are relatively dense and structured. In contrast, point clouds captured in outdoor environments, particularly used in autonomous driving, are significantly sparser and often acquired under harsh conditions, including occlusions and uneven sampling. To show MSCN’s effectiveness in these challenges, we extend our experiments to AV datasets.
V-D Evaluation on AV Datasets: Robustness to Different Sensor Configurations and Geometric Deformations
@@
Our methods focus on maintaining consistent object recognition performance regardless of the type of point cloud data encountered. This means MSCN was designed to work well not only with the data they were originally trained on but also with data from different sources. To thoroughly assess their effectiveness in the AV dataset, we conduct two types of evaluations. Both intra-domain and cross-domain performances are evaluated to provide a comprehensive understanding of the capabilities of the MSCN and to ensure its reliability and effectiveness across various AV datasets.
@@
Table IV shows the classification performance within a single domain. These results measure the model’s ability to classify point clouds within the same domain used for training. While MSCN achieves competitive results, it lags slightly behind the SOTA methods in terms of accuracy. This performance gap can be explained from the bias-variance trade-off [49]. MSCN, designed for robust domain-invariant feature extraction, prioritizes reducing model variance to generalize effectively across diverse domains. As a result, its intra-domain performance exhibits slightly higher bias, which is a common trade-off observed in models designed with high generalization capacity. Specifically, while other SOTA models may minimize bias within the training domain, they often exhibit poor performance in domain-shift scenarios due to higher variance.
To investigate SOTA models’ sensitivity and MSCN’s robustness under the different LiDAR configurations, we further evaluate its performance under cross-domain scenarios as illustrated in Table V, where substantial domain shifts between training and testing datasets. PointNet, PointNeXt, PointMLP, DGCNN, PointTransformer, and PointMamba demonstrate good performance when the domain shifts from simulation to the real world (Sim2Real) but show significant performance degradation when shifting from one real-world dataset to another (Real2Real). This is because real-world datasets not only differ in resolution but also exhibit harsh characteristics such as irregular sampling, occlusion, and sensor noise. These models tend to overfit to specific point cloud properties, leading to substantial recognition failures when deployed across diverse real-world LiDARs, which poses a critical risk for safety-critical applications in AVs. On the other hand, GCN and SCN show robust performance in Real2Real domain change scenarios but are sensitive in Sim2Real scenarios. This sensitivity arises because their graph-based structures adapt well to irregular real-world point clouds but struggle to generalize from the uniform and noise-free distribution of simulated point clouds. Consequently, their limited ability to bridge the Sim2Real gap hinders the use of large-scale simulation data for training, thereby restricting scalability and reducing the cost-effectiveness of model development.
@@
However, our MSCN demonstrates robust performance in both Real2Real and Sim2Real scenarios compared to the aforementioned methods. In particular, MSCN surpasses all SOTA methods with an average cross-domain accuracy of 82%, significantly higher than PointTransformer’s 66.2%. This 15.8% improvement highlights the effectiveness of MSCN in handling domain shifts, which is critical for the deployment of AVs.
@@
In particular, the average inference time of MSCN is only 3.8 ms, which is well suited for AVs. Modern rotary LiDAR sensors typically operate at speeds of 10 Hz to 20 Hz, meaning one full rotation takes 0.05 to 0.1 seconds. Given that MSCN can process a single point cloud within 3.8 ms, it is capable of handling multiple scans within a single rotation, ensuring real-time performance without bottlenecking the data pipeline. This efficient processing time, combined with robust cross-domain performance, demonstrates the viability of MSCN for deployment in AV systems.
TABLE V: Cross-domain classification across both real-world AV and simulation datasets.
KITTI (64CH) [36] PanKyo (128CH) [41] nuScenes (32CH) [37] Simulation (32CH) [42] Average
Methods Param. (M) PanKyo nuScenes KITTI nuScenes KITTI PanKyo KITTI PanKyo nuScenes Acc(%)
PointNet [12] 3.0 59.3 62.8 6.2 21.8 36.3 47.9 82.6 78.5 71.8 51.9
PointNeXt [40] 14.1 67.6 31.4 5.8 17.3 31.6 18.6 82.7 88.6 76.1 46.6
PointMLP [15] 13.2 17.7 80.4 87.7 64 47.1 15.2 81 84.6 35.2 57
DGCNN [14] 4.2 14.5 38 6.2 25.6 72.2 20.1 81.2 82.7 94.6 48.3
GCN [16] 1.2 60.9 60.6 83 64.1 92.8 86.2 20.7 16.7 23.8 56.5
SCN [41] 3.5 42.6 72.3 84.4 67.3 92.1 74.3 30.9 48.4 29.1 60.2
PointTransformer [47] 9.6 7.9 88.1 77.3 73.3 79.1 8.8 80.5 87.1 94.0 66.2
PointMamba[46] 12.3 6.6 90.1 11.5 37.5 61.6 30.6 96.9 90.9 89.3 57.2
MSCN (Ours) 4.2 81.7 75.3 88.2 78.5 92.8 84.8 83.5 83.9 69.2 82
KITTI [36] PanKyo [41] nuScenes [37] Simulation [42] Avg.
Methods Param. KITTI PanKyo nuScenes KITTI PanKyo nuScenes KITTI PanKyo nuScenes KITTI PanKyo nuScenes Time
PointNet [12] 3.0 4.1 2.0 0.8 3.9 1.9 0.7 3.9 2.2 0.8 4.0 2.1 0.8 2.3
PointNeXt [40] 14.1 38.2 37.5 37.3 38.2 37.4 37.1 38.1 37.5 37.2 38.1 37.4 37.2 37.6
PointMLP [15] 13.2 32.4 31.1 31.0 32.4 31.1 31.0 32.6 31.4 31.1 32.5 31.2 31.1 31.6
DGCNN [14] 4.2 10 8.8 8.2 10 8.8 8.2 10 8.8 8.1 10.1 8.9 8.2 9.0
GCN [16] 1.2 3.6 2.9 2.6 3.6 3.0 2.6 3.6 2.9 2.6 3.6 2.8 2.5 3.0
SCN [41] 3.5 4.0 3.2 2.8 4.2 3.4 2.8 4.5 3.3 2.9 4.1 3.1 2.7 3.4
PointTransformer [47] 9.6 36.5 34.6 34.1 35.8 36.3 35.3 36.5 34.6 34.1 36.4 34.5 34.0 35.2
PointMamba[46] 12.3 5.4 4.1 3.4 5.4 4.1 3.4 5.5 4.1 3.4 5.4 4.1 3.4 4.3
MSCN (Ours) 4.2 4.7 3.6 3.1 4.9 3.6 3.2 4.4 3.6 3.2 4.7 3.6 3.1 3.8
Figure 6: Robustness evaluation under rotations and shifts in both intra- and cross-domain settings.
TABLE VII: Layer configuration evaluation results on both real-world AV and simulation datasets.
@@
KITTI [36] PanKyo only SCL 78.9
@@
PanKyo [41] KITTI only SCL 84.9
@@
nuScenes [37] KITTI only SCL 92.1
@@
Simulation [42] PanKyo only SCL 75.1
@@
TABLE VIII: Unseen domain generation effects on both real-world AV and simulation datasets.
@@
KITTI [36] PanKyo MSCN 81.7 3.6
@@
Pankyo [41] PanKyo MSCN 94.5 3.6
@@
nuScenes [37] PanKyo MSCN 84.8 3.6
@@
Simulation [42] PanKyo MSCN 83.9 3.6
@@ V-E Ablation study
We evaluate the compatibility of layer configurations of MSCN and the performance enhancement when MSCN was trained with synthetic point cloud datasets from our unseen domain generation method.
@@
To evaluate the impact of our unseen domain generalization methods on improving the leaning of domain-invariant feature representation, we perform various experiments based on both intra-domain and cross-domain classification, which are illustrated in Table VIII. MSCN, which can increase the amount of training without overfitting, performs better than MSCN for single-domain classification. Note that MSCN does not include a domain generation module. With the domain generation module, MSCN demonstrates even more robust performance than MSCN in various domain change scenarios, with an average improvement of 5.8%, indicating that our proposed unseen domain generation method ensures domain generalization ability.
@@ VI CONCLUSION
In this paper, we introduced MSCN, which learns domain-invariant feature representations by capturing the geometric information of 3D point clouds from the standpoint of both local context and overall context. Moreover, we developed the MSCN through our unseen domain expansion framework, enabling the MSCN to learn more robust feature representations. Our MSCN and MSCN demonstrated the highest performance in cross-domain classification tasks across all AV datasets, which are derived not only from real-world data, different regions, and sensors, but also from simulators. These results indicate that our method addresses sensor-dependent problems in point cloud classification tasks for the deployment of AVs. In light of these findings, our future research endeavors will focus on expanding our method to object detection tasks for AV systems that require domain-invariant recognition.
@@ Acknowledgment
This work was supported by the Korea National Police Agency(KNPA) under the project ‘Development of the autonomous driving patrol service for active prevention and response to traffic accidents’ (RS-2024-00403630).

### v5

@@ V-C Evaluation on Indoor Real-World Datasets: Comparable Performance and Robustness
MSCN(ours) 92.0
V-D Evaluation on Self-Driving Datasets: Cross-Domain Robustness
@@
Our methods focus on maintaining consistent object recognition performance regardless of the type of point cloud data encountered. This means they are designed to work well not only with the data they were originally trained on but also with data from different sources. To thoroughly assess their effectiveness, we conducted two types of evaluations. Both intra-domain and cross-domain performances were evaluated to provide a comprehensive understanding of the capabilities of the methods and to ensure their reliability and effectiveness across various datasets.
@@
First, the performances are demonstrated within the KITTI [37], nuScenes [38], and PanKyo [42] datasets to capture classification performance seamlessly using the training dataset on the test dataset. Table IV shows the classification performance within a single domain. These results measure the model’s ability to classify point clouds within the same domain used for training. While MSCN achieves competitive results, it lags slightly behind the SOTA methods in terms of accuracy. For example, the average intra-domain accuracy of MSCN across all datasets is 94.7%, compared to PointTransformer’s 98%. This performance gap can be explained from the bias-variance trade-off [50, 51]. MSCN, designed for robust domain-invariant feature extraction, prioritizes reducing model variance to generalize effectively across diverse domains. As a result, its intra-domain performance exhibits slightly higher bias, which is a common trade-off observed in models designed with high generalization capacity. Specifically, while other SOTA models may minimize bias within the training domain, they often exhibit poor performance in domain-shift scenarios due to higher variance.
To demonstrate MSCN’s robustness, we further evaluated its performance under cross-domain scenarios as illustrated in Table V, where substantial domain shifts between training and testing datasets. PointNet, PointNeXt, PointMLP, DGCNN, PointTransformer, and PointMamba demonstrate good performance when the domain shifts from simulation to the real world (Sim2Real) but show significantly lower performance when shifting from one real-world dataset to another (Real2Real). On the other hand, GCN and SCN show robust performance in Real2Real domain change scenarios but are sensitive in Sim2Real scenarios.
@@
However, our MSCN demonstrates robust performance in both Real2Real and Sim2Real scenarios compared to the aforementioned methods. In particular, MSCN surpasses all SOTA methods with an average cross-domain accuracy of 82%, significantly higher than PointTransformer’s 66.2%. This 15.8% improvement highlights the effectiveness of MSCN in handling domain shifts, which is critical for real-world applications such as autonomous driving and robotics.
@@
TABLE V: Cross-domain classification results on KITTI [37], nuScenes [38], and PanKyo [42] datasets. The bold black, blue indicate the first and the second best performance on each scenario, respectively.
[…]
! KITTI (64CH) PanKyo (128CH) nuScenes (32CH) Simulation (32CH) Average Methods Param. (M) PanKyo nuScenes KITTI nuScenes KITTI PanKyo KITTI PanKyo nuScenes Acc(%) PointNet [11] 3.0 59.3 62.8 6.2 21.8 36.3 47.9 82.6 78.5 71.8 51.9 PointNeXt [41] 14.1 67.6 31.4 5.8 17.3 31.6 18.6 82.7 88.6 76.1 46.6 PointMLP [14] 13.2 17.7 80.4 87.7 64.0 47.1 15.2 81.0 84.6 35.2 57.0 DGCNN [13] 4.2 14.5 38.0 6.2 25.6 72.2 20.1 81.2 82.7 94.6 48.3 GCN [15] 1.2 60.9 60.6 83.0 64.1 92.8 86.2 20.7 16.7 23.8 56.5 SCN [42] 3.5 42.6 72.3 84.4 67.3 92.1 74.3 30.9 48.4 29.1 60.2 PointTransformer [48] 9.6 7.9 88.1 77.3 73.3 79.1 8.8 80.5 87.1 94.0 66.2 PointMamba [47] 12.3 6.6 90.1 11.5 37.5 61.6 30.6 96.9 90.9 89.3 57.2 MSCN(ours) 4.2 81.7 75.3 88.2 78.5 92.8 84.8 83.5 83.9 69.2 82.0
[…]
! KITTI PanKyo nuScenes Simulation Average Methods Param.(M) KITTI PanKyo nuScenes KITTI PanKyo nuScenes KITTI PanKyo nuScenes KITTI PanKyo nuScenes Time(ms) PointNet [11] 3.0 4.1 2.0 0.8 3.9 1.9 0.7 3.9 2.2 0.8 4.0 2.1 0.8 2.3 PointNeXt [41] 14.1 38.2 37.5 37.3 38.2 37.4 37.1 38.1 37.5 37.2 38.1 37.4 37.2 37.6 PointMLP [14] 13.2 32.4 31.1 31.0 32.4 31.1 31.0 32.6 31.4 31.1 32.5 31.2 31.1 31.6 DGCNN [13] 4.2 10.0 8.8 8.2 10.0 8.8 8.2 10.0 8.8 8.1 10.1 8.9 8.2 9.0 GCN [15] 1.2 3.6 2.9 2.6 3.6 3.0 2.6 3.6 2.9 2.6 3.6 2.8 2.5 3.0 SCN [42] 3.5 4.0 3.2 2.8 4.2 3.4 2.8 4.5 3.3 2.9 4.1 3.1 2.7 3.4 PointTransformer [48] 9.6 36.5 34.6 34.1 35.8 36.3 35.3 36.5 34.6 34.1 36.4 34.5 34.0 35.2 PointMamba [47] 12.3 5.4 4.1 3.4 5.4 4.1 3.4 5.5 4.1 3.4 5.4 4.1 3.4 4.3 MSCN(ours) 4.2 4.7 3.6 3.1 4.9 3.6 3.2 4.4 3.6 3.2 4.7 3.6 3.1 3.8
[…]
Figure 6: Evaluation of Invariance Properties. All models were trained on the PanKyo [42] dataset and then tested on (a) PanKyo[42], (b) KITTI[37], and (c) nuScenes[38] datasets. The graphs on the left illustrate the performance of the models when point clouds were rotated along the upward direction. The graphs on the right show the results when the models were tested with randomly shifted point clouds within a certain distance in all directions.
@@
In particular, the average inference time of MSCN is only 3.8 ms, which is well suited for real-world applications such as autonomous driving. Modern rotary LiDAR sensors typically operate at speeds of 10 Hz to 20 Hz, meaning one full rotation takes 0.05 to 0.1 seconds. Given that MSCN can process a single point cloud within 3.8 ms, it is capable of handling multiple scans within a single rotation, ensuring real-time performance without bottlenecking the data pipeline. This efficient processing time, combined with robust cross-domain performance, demonstrates the viability of MSCN for deployment in real-world autonomous systems.
TABLE VII: Layer configuration evaluation results on KITTI [37], nuScenes [38], PanKyo [42], and simulation [43] datasets.
@@
KITTI PanKyo only SCL 78.9
@@
PanKyo KITTI only SCL 84.9
@@
nuScenes KITTI only SCL 92.1
@@
Simulation PanKyo only SCL 75.1
@@
TABLE VIII: Unseen domain generation effects on KITTI [37], nuScenes [38], PanKyo [42], and simulation [43] datasets. The $\dagger$} mark indicates the method trained with our unseen domain generation method, achieved from equation 13 and equation 19.
@@
KITTI PanKyo MSCN 81.7 3.6
@@
Pankyo PanKyo MSCN 94.5 3.6
@@
nuScenes PanKyo MSCN 84.8 3.6
@@
Simulation PanKyo MSCN 83.9 3.6
@@
TABLE IX: Loss ablation results of MSCN trained on KITTI dataset and evaluated across KITTI [37], nuScenes [38], and PanKyo [42].
Training Dataset Test Dataset Method Acc.(%)
KITTI KITTI w/o $L_{recon}$ 99.7
w/o $L_{div}$ 99.6
w/o $L_{adv}$ 99.4
w/o $L_{gen,ce}$ 99.7
MSCN 99.7
PanKyo w/o $L_{recon}$ 78.4
w/o $L_{div}$ 71.9
w/o $L_{adv}$ 59.8
w/o $L_{gen,ce}$ 63.2
MSCN 81.7
nuScenes w/o $L_{recon}$ 82.1
w/o $L_{div}$ 80.4
w/o $L_{adv}$ 79.0
w/o $L_{gen,ce}$ 81.6
MSCN 82.7
@@ V-E Ablation study
We evaluate the compatibility of layer configurations of MSCN, the performance enhancement when MSCN was trained with synthetic point cloud datasets from our unseen domain generation method, and the impact of each loss function used for domain expansion.
@@
To evaluate the impact of our unseen domain generalization methods on improving the leaning of domain-invariant feature representation, we performed various experiments based on both intra-domain and cross-domain classification, which are illustrated in Table VIII. MSCN, which can increase the amount of training without overfitting, performs better than MSCN for single-domain classification. Note that MSCN does not include a domain generation module. With the domain generation module, MSCN demonstrates even more robust performance than MSCN in various domain change scenarios, with an average improvement of 5.8%, indicating that our proposed unseen domain generation method ensures domain generalization ability.
[…]
@@ VI Conclusion
In this paper, we introduced MSCN, which learns domain-invariant feature representations by capturing the geometric information of 3D point clouds from the standpoint of both local context and overall context. Furthermore, we developed an unseen-domain expansion framework that enables MSCN to learn more robust and transferable features across different sensing conditions. Comprehensive cross-domain evaluations demonstrated that our MSCN achieved an average accuracy of 82.0%, surpassing the strong baseline PointTransformer by 15.8%, while MSCN further improved performance by an additional 5.8%. These quantitative results confirm the effectiveness of our approach in addressing sensor-dependent and domain-shift challenges in point cloud classification for autonomous vehicles. In light of these findings, our future research endeavors will focus on expanding our method to object detection tasks for real-world applications that require domain-invariant recognition, such as autonomous driving, robotics, and virtual reality.
@@ Acknowledgment
This work was supported by the Korea National Police Agency(KNPA) under the project ‘Development of the autonomous driving patrol service for active prevention and response to traffic accidents’ (RS-2024-00403630) and the BK21 FOUR program of the National Research Foundation Korea funded by the Ministry of Education(NRF5199991014091).